# Dynamics of Pseudoentanglement

### v1

@@ D.1 Single particle
In this Appendix we review, for completeness, the solution of the simple random walk on the hypercube $\mathbb{Z}_{2}^{N}$ Diaconis et al. (1990). We label each vertex of the hypercube by a bitstring $\mathbf{z}\in\{0,1\}^{N}$; two vertices $\mathbf{x},\mathbf{y}$ are neighbors iff the corresponding bitstrings have Hamming distance 1: $|\mathbf{x}\oplus\mathbf{y}^{\prime}|=1$ ($\oplus$ is sum modulo 2). In each time step the walker hops randomly in any one of the $N$ directions, represented by the Markov chain transition matrix
@@
Note that as written, the walk is not aperiodic: the parity of the Hamming weight alternates in each time step (i.e., the hypercube is bipartite). This problem can be eliminated by allowing idling, i.e. replacing $\Gamma_{\mathbf{x},y}\mapsto p\Gamma_{\mathbf{x},y}+(1-p)\delta_{\mathbf{x},\mathbf{y}}$ (the walker takes a step with probability $p$, remains in place otherwise). This does not change the eigenfunctions and changes the eigenvalues in a trivial way, $\lambda\mapsto p\lambda+(1-p)$.
@@
Here we have introduced $\mathbf{r}\equiv\mathbf{x}\oplus\mathbf{y}$ and used $\sum_{\mathbf{x}}(-1)^{\mathbf{x}\cdot\mathbf{c}}=D\delta_{\mathbf{c},\boldsymbol{0}}$. The eigenvalue associated to “wave vector” $\mathbf{a}\in\{0,1\}^{N}$ is
@@
There is a non-degenerate $\lambda_{0}=1$ eigenvalue, associated to $\mathbf{a}=\boldsymbol{0}$ (the uniform state), then a gap of magnitude $2/N$, and then an $N$-fold degenerate eigenvalue $\lambda_{1}=1-2/N$. The relaxation time is therefore $t_{\rm rel}=1/(1-\lambda_{1})=N/2$. One can also identify the mixing time exactly and prove that the problem (with idling) features a cutoff Diaconis et al. (1990). A simple way to estimate the mixing time is to require that $N\lambda_{1}^{t}$ be small, which gives $t_{\rm mix}\sim N\log(N)$ (the correct scaling up to constant factors).
The gate set of Sec. IV.3 induces a simple random walk on the hypercube with idling parameter $p=1/4$, giving rescaled eigenvalues $\lambda_{1}=1-\frac{2}{N}\mapsto\frac{1}{4}\lambda_{1}+\frac{3}{4}=1-\frac{1}{2N}$, thus the relaxation time $t_{\rm rel}=2N$.
@@ D.2 Many particles: multipole eigenmodes
where for each bitstring $\mathbf{a}\in\mathbb{Z}_{2}^{N}$ we defined $Z_{\mathbf{a}}=\bigotimes_{i=1}^{N}Z_{i}^{a_{i}}$, and $n_{S}(Z_{\mathbf{a}}=\pm 1)$ is the number of elements $\mathbf{z}\in S$ such that $\bra{\mathbf{z}}Z_{\mathbf{a}}\ket{\mathbf{z}}=\pm 1$. The action of the transition matrix $\Gamma_{S,S^{\prime}}$ on these functions is:
$\sum_{S^{\prime}}\Gamma_{S,S^{\prime}}f_{\mathbf{a}}(S^{\prime})$ $=\sum_{S^{\prime}}\frac{|\{u\in\mathcal{G}:uS=S^{\prime}\}|}{|\mathcal{G}|}f_{\mathbf{a}}(S^{\prime})$
$=\sum_{u\in\mathcal{G}}\frac{1}{|\mathcal{G}|}f_{\mathbf{a}}(uS)$ (94)
Here we have used the definition of the transition probability $\Gamma_{S,S^{\prime}}$ as the fraction of gates $u$ in the gate set $\mathcal{G}$ that maps subset $S$ to $S^{\prime}$, and then changed variables via $S^{\prime}\equiv uS$. Summing over $u$ over-counts each subset $S^{\prime}$ in the original sum by a factor of |$|\{u^{\prime}\in\mathcal{G}:u^{\prime}S=uS\}|$, which precisely cancels the numerator in $\Gamma_{S,S^{\prime}}$. Now we note that
$f_{\mathbf{a}}(uS)\propto 1-2n_{uS}(Z_{\mathbf{a}}=-1)=1-2n_{S}(uZ_{\mathbf{a}}u=-1).$ (95)
If $u$ is the $\textsf{CC}X$ gate that flips qubit $i$ controlled by its neighbors $i\pm 1$ being in states $c_{\pm}\in\{-1,1\}$, we have
$uZ_{i}u=\left[I-2\frac{I+c_{+}Z_{i+1}}{2}\frac{I+c_{-}Z_{i-1}}{2}\right]Z_{i}$ (96)
and $uZ_{j}u=Z_{j}$ for all $j\neq i$. The operator $uZ_{i}u$ has eigenvalues ±$\pm 1$ in equal numbers. The $+1$ eigenspace is specified by $Z_{i}=-1$ and $c_{-}Z_{i-1}=c_{+}Z_{i+1}=+1$ ($D/8$ states), or $Z_{i}=+1$ and $c_{\xi}Z_{i+\xi}=-1$ for at least one $\xi\in\{-1,+1\}$ ($83D/$ 8 states). The remaining $D/2$ computational basis states form the $-1$ eigenspace. It is easy to see that each computational basis state with $Z_{i}=+1$ is in the $+1$ eigenspace of $uZ_{i}u$ for 3 choices of control bits, and in the $-1$ eigenspace for one choice ($c_{+}=c_{-}=+1$). Therefore, when we sum over the control bits $c_{\pm}\in\{0,1\}$ (as part of our sum over $u\in\mathcal{G}$) we find $\sum_{c_{\pm}}n_{S}(u_{i,c_{\pm}}Z_{\mathbf{a}}u_{i,c_{\pm}}=+1)=(4-2\delta_{a_{i},1})n_{S}(Z_{i}=+1)$. On net, we find
$\frac{1}{4N}\sum_{u\in\mathcal{G}}f_{\mathbf{a}}(uS)$ $=\frac{1}{4N}\sum_{i=1}^{N}\sum_{c_{\pm}=\pm 1}f_{\mathbf{a}}(u_{i,c_{-},c_{+}}S)$
@@
$=\left(1-\frac{|\mathbf{a}|}{2N}\right)f_{\mathbf{a}}(S)=\lambda_{\mathbf{a}}f_{\mathbf{a}}(S),$ (97)
@@
$\overline{\langle Z_{\mathbf{a}}(t)\rangle}$ $=\sum_{S}p_{t}(S)\bra{\psi_{S}}Z_{\mathbf{a}}\ket{\psi_{S}}=(p_{t}|f_{\mathbf{a}})$
$=(p_{0}|\Gamma^{t}|f_{\mathbf{a}})=\left(1-\frac{|\mathbf{a}|}{2N}\right)^{t}(p_{0}|f_{\mathbf{a}}),$ (98)
@@ D.3 Two particles: relative coordinate eigenmodes
Here we provide more details on eigenmodes of the two-particle problem that depend subset $S=\{\mathbf{z},\mathbf{z}^{\prime}\}$ only through the relative coordinate $\mathbf{r}=\mathbf{z}\oplus\mathbf{z}^{\prime}$. Let us consider how $\mathbf{r}$ might change under the action of our local automaton gate set $\mathcal{G}=\{u_{iab}\}$. There are two cases to consider:
@@
The rules above describe a random walk on $\mathbb{Z}_{2}^{N}$ based on “activated” bit flips, where $r_{i}$ can flip only if one of its neighbors $r_{i\pm 1}$ is 1. Clearly $\mathbf{r}=\boldsymbol{0}$ is a steady state, but it does not describe a valid subset (it gives $\mathbf{z}=\mathbf{z}^{\prime}$). One can check that the restriction to $\mathbb{Z}_{2}^{N}\setminus\{\boldsymbol{0}\}$ is irreducible, aperiodic (since lazy moves are allowed) and reversible, thus it has the uniform distribution as unique steady state.
@@
The linear-in-$N$ relaxation time corresponds to the Hamiltonian $\mathcal{H}$ being gapped.
@@ Appendix E Properties of the local automaton gate set
In this Appendix we give more details on the local automaton gate set $\mathcal{G}$ used in Sec. IV.3. We recall that the gate set is defined as $\mathcal{G}=\{u_{iab}=\textsf{C}_{i-1,a}\textsf{C}_{i+1,b}{X}_{i}\}$, with $i\in[N]$, $a,b\in\{0,1\}$, and periodic boundary conditions (i.e., $i\pm 1$ are taken modulo $N$). Each time step consists of drawing a gate $u_{iab}\in\mathcal{G}$ uniformly at random and applying it to the system. This defines a random walk on the permutation group $\mathcal{S}_{D}$ with transition matrix
@@
(we use $\Gamma$ for all these transition matrices and rely on the indices to distinguish them).
@@
It is well known Shende et al. (2003); Aaronson et al. (2015) that the Toffoli ($\textsf{CC}X$) and NOT ($X$) gates are “almost” universal for classical computation, in the sense that they generate the subgroup of $\mathcal{S}_{D}$ made of even permutations (permutations that can be decomposed into an even number of transpositions). Our gate set includes Toffoli gates between consecutive qubits ($u_{i11}$); further, it can generate NOT gates ($X_{i}=u_{i00}u_{i01}u_{i10}u_{i11}$) and CNOT gates between neighboring qubits, with either choice of control and target:
@@
Finally, by composing local SWAPs we can make arbitrary permutations of the qubits, and thus generate the whole (non-local) Toffoli + NOT gate set. Therefore $\mathcal{G}$ generates all even permutations. Now, given two sets $S,S^{\prime}\in\Sigma_{m}$, we can find a permutation $\sigma\in\mathcal{S}_{D}$ that maps one to the other: $S^{\prime}=\sigma(S)$. If $\sigma$ is even, we are done. If $\sigma$ is odd, then let $i,j\in[D]$ be two basis states that are not in subset $S$ (we have assumed $|S|=m\leq D-2$ so that two such states exist), and let $\tau_{ij}\in\mathcal{S}_{D}$ be the transposition $i\leftrightarrow j$. Then $u\circ\tau_{ij}$ is an even permutation that maps $S$ to $S^{\prime}$. Since any two sets can be connected by an even permutation and any even permutation can be decomposed into a product of $u\in\mathcal{G}$, the Markov chain can connect any two subsets $S,S^{\prime}\in\Sigma_{m}$ and is thus irreducible. ∎
@@
Reversibility, irreducibility and aperiodicity imply that the Markov chain has the uniform distribution $\pi_{m}$ as its unique steady state Levin and Peres (2017). This proves that our 3-qubit automaton gate set equilibrates to the uniform distribution on all subset states with $|S|=m\leq D/4$.
Furthermore, it is easy to see that three-qubit gates are necessary to obtain the above result. Indeed two-qubit automaton gates (generated by NOT and CNOT) are affine transformations, in the sense that for any such gate $u$ we have $u\ket{\mathbf{z}}=\ket{A_{u}\mathbf{z}\oplus\mathbf{b}_{u}}$, with $A_{u}$ a binary matrix and $\mathbf{b}_{u}$ a bitstring. Any transformations that are not affine cannot be generated by two-qubit automaton gates. As it turns out, the gate set comprising NOT and CNOT gates satisfies irreducibility up to $m=3$; an example of two subsets that cannot be connected with $m=4$ and $N=3$ is given by
@@
(each row is a bitstring $\mathbf{z}$ in the set). It is easy to see that the sum (modulo 2) of bitstrings in each set, $\mathbf{z}_{\rm tot}$, transforms under automaton gate $u$ according to $\mathbf{z}_{\rm tot}\mapsto A_{u}\mathbf{z}_{\rm tot}$ (the additive term $\mathbf{b}_{u}$ appears 4 times and thus cancels). But $S$ has $\mathbf{z}_{\rm tot}=(0,0,0)$ while $S^{\prime}$ has $\mathbf{z}_{\rm tot}=(1,1,1)$, so any sequence of $u$’s mapping $S$ to $S^{\prime}$ must be such that $\prod_{u}A_{u}$ annihilates $(1,1,1)$. This is incompatible with a reversible computation (e.g. it would imply a collision between strings $000$ and $111$).
Another way of arriving at a similar conclusion is that pseudoentanglement requires magic Grewal et al. (2023), and two-qubit automaton gate sets happen to also be Clifford. Thus if we start from a stabilizer subset state (e.g. |$\ket{0}^{\otimes N-k}\otimes\ket{+}^{\otimes k}$) we cannot produce a pseudoentangled ensemble. However, for $m\leq 3$, it is instead possible to show that the gate set comprising NOT and CNOT yields irreducible Markov chains. This is perhaps related to the fact that mimicking the 3 lowest moments of the Haar measure does not require magic (as Clifford circuits form 3-designs on qubits).

### v3

@@ D.1 Single particle
In this Appendix we review, for completeness, the solution of the simple random walk on the hypercube $\mathbb{Z}_{2}^{N}$ diaconis_asymptotic_1990 . We label each vertex of the hypercube by a bitstring $\mathbf{z}\in\{0,1\}^{N}$; two vertices $\mathbf{x},\mathbf{y}$ are neighbors iff the corresponding bitstrings have Hamming distance 1: $|\mathbf{x}\oplus\mathbf{y}^{\prime}|=1$ ($\oplus$ is sum modulo 2). In each time step the walker hops randomly in any one of the $N$ directions, represented by the Markov chain transition matrix
@@
Note that as written, the walk is not aperiodic: the parity of the Hamming weight alternates in each time step (this is a general feature of bipartite graphs). This problem can be eliminated by allowing idling, i.e. replacing $\Gamma_{\mathbf{x},y}\mapsto p\Gamma_{\mathbf{x},y}+(1-p)\delta_{\mathbf{x},\mathbf{y}}$ (the walker takes a step with probability $p$, remains in place otherwise). This does not change the eigenfunctions and changes the eigenvalues in a trivial way, $\lambda\mapsto p\lambda+(1-p)$.
@@
Here we have introduced $\mathbf{r}\equiv\mathbf{x}\oplus\mathbf{y}$ and used $\sum_{\mathbf{x}}(-1)^{\mathbf{x}\cdot\mathbf{c}}=D\delta_{\mathbf{c},\bm{0}}$. The eigenvalue associated to “wave vector” $\mathbf{a}\in\{0,1\}^{N}$ is
@@
There is a non-degenerate $\lambda_{0}=1$ eigenvalue, associated to $\mathbf{a}=\bm{0}$ (the uniform state), then a gap of magnitude $2/N$, and then an $N$-fold degenerate eigenvalue $\lambda_{1}=1-2/N$. The relaxation time is therefore $t_{\rm rel}=1/(1-\lambda_{1})=N/2$. One can also identify the mixing time exactly and prove that the problem (with idling) features a cutoff diaconis_asymptotic_1990 . A simple way to estimate the mixing time is to require that $N\lambda_{1}^{t}$ be small, which gives $t_{\rm mix}\sim N\log(N)$ (the correct scaling up to constant factors).
The all-to-all gate set of Sec. 4.2 induces a simple random walk on the hypercube with idling parameter $p=2^{1-N}=2/D$, giving rescaled eigenvalues $\lambda_{1}=1-\frac{2}{N}\mapsto\frac{2}{D}\lambda_{1}+1-\frac{2}{D}=1-\frac{4}{DN}$, hence the relaxation time $t_{\rm rel}=ND/4$. Similarly the local gate set of Sec. 4.3 induces a simple random walk on the hypercube with idling parameter $p=1/4$, giving $\lambda_{1}=1-\frac{2}{N}\mapsto\frac{1}{4}\lambda_{1}+\frac{3}{4}=1-\frac{1}{2N}$, hence the relaxation time $t_{\rm rel}=2N$.
@@ D.2 Many particles: multipole eigenmodes
where for each bitstring $\mathbf{a}\in\{0,1\}^{N}$ we defined $Z_{\mathbf{a}}=\bigotimes_{i=1}^{N}Z_{i}^{a_{i}}$, and $n_{S}(Z_{\mathbf{a}}=\pm 1)$ is the number of elements $\mathbf{z}\in S$ such that $\bra{\mathbf{z}}Z_{\mathbf{a}}\ket{\mathbf{z}}=\pm 1$. The action of the transition matrix $\Gamma_{S,S^{\prime}}$ on these functions is:
$\sum_{S^{\prime}}\Gamma_{S,S^{\prime}}f_{\mathbf{a}}(S^{\prime})$ $=\sum_{S^{\prime}}\frac{|\{u\in\mathcal{G}:u(S)=S^{\prime}\}|}{|\mathcal{G}|}f_{\mathbf{a}}(S^{\prime})$
$=\sum_{u\in\mathcal{G}}\frac{1}{|\mathcal{G}|}f_{\mathbf{a}}(u(S))$ (77)
Here we have used the definition of the transition probability $\Gamma_{S,S^{\prime}}$ as the fraction of gates $u$ in the gate set $\mathcal{G}$ that maps subset $S$ to $S^{\prime}$, and then changed variables via $S^{\prime}\equiv u(S)$. Summing over $u$ over-counts each subset $S^{\prime}$ in the original sum by a factor of |$|\{u^{\prime}\in\mathcal{G}:u^{\prime}(S)=u(S)\}|$, which precisely cancels the numerator in $\Gamma_{S,S^{\prime}}$. Now we note that
$f_{\mathbf{a}}(u(S))=\frac{2}{m}n_{u(S)}(Z_{\mathbf{a}}=1)-1=\frac{2}{m}n_{S}(uZ_{\mathbf{a}}u=1)-1.$ (78)
Let $u\equiv u_{i,c_{-},c_{+}}$ ($i$ is the location of the target, $c_{\pm}$ are the control bits for sites $i\pm 1$). If $a_{i}=0$, then $uZ_{\mathbf{a}}u=u^{2}Z_{\mathbf{a}}=Z_{\mathbf{a}}$. If $a_{i}=1$, we define $Z_{\mathbf{a}}\equiv Z_{\mathbf{b}}Z_{i}$, so that $uZ_{\mathbf{a}}u=Z_{\mathbf{b}}uZ_{i}u$. The operator $uZ_{i}u$ is still diagonal in the computational basis. A bitstring has $uZ_{\mathbf{a}}u=+1$ if and only if $Z_{\mathbf{b}}=uZ_{i}u$. Any given bitstring $\mathbf{z}\in S$ will change its value of $Z_{i}$ under gate $u_{i,c_{-},c_{+}}$ for exactly one out of four possible choices for the control bits $c_{\pm}$ (namely $c_{\pm}=z_{i\pm 1}$). It follows that
$\sum_{c_{\pm}=0,1}n_{S}(Z_{\mathbf{b}}=u_{i,c_{-},c_{+}}Z_{i}u_{i,c_{-},c_{+}}=1)=3n_{S}(Z_{\mathbf{b}}=Z_{i}=1)+n_{S}(Z_{\mathbf{b}}=-Z_{i}=1)$ (79)
and
$\sum_{c_{\pm}=0,1}n_{S}(Z_{\mathbf{b}}=u_{i,c_{-},c_{+}}Z_{i}u_{i,c_{-},c_{+}}=-1)=3n_{S}(Z_{\mathbf{b}}=Z_{i}=-1)+n_{S}(Z_{\mathbf{b}}=-Z_{i}=-1).$ (80)
Adding the two gives
$\sum_{c_{\pm}=0,1}n_{S}(u_{i,c_{-},c_{+}}Z_{\mathbf{a}}u_{i,c_{-},c_{+}}=1)$ $=3n_{S}(Z_{\mathbf{a}}=1)+n_{S}(Z_{\mathbf{a}}=-1)$
$=m+2n_{S}(Z_{\mathbf{a}}=1),$ (81)
where we used the fact that $n_{S}(Z_{\mathbf{a}}=1)+n_{S}(Z_{\mathbf{a}}=-1)=|S|=m$. Therefore, when we sum over the control bits $c_{\pm}\in\{0,1\}$ as part of our sum over $u\in\mathcal{G}$ in Eq. (77), if $a_{i}=1$ we find
$\sum_{c_{\pm}=\pm 1}f_{\mathbf{a}}(u_{i,c_{-},c_{+}}(S))$ $=\frac{2}{m}[m+2n_{S}(Z_{\mathbf{a}}=1)]-4=2f_{\mathbf{a}}(S)$ (82)
Combining this with the result for $a_{i}=0$ gives
$\sum_{c_{\pm}=\pm 1}f_{\mathbf{a}}(u_{i,c_{-},c_{+}}(S))=2(2-\delta_{a_{i},1})f_{\mathbf{a}}(S).$ (83)
Finally, plugging this into Eq. (77), we obtain
$\frac{1}{4N}\sum_{u\in\mathcal{G}}f_{\mathbf{a}}(u(S))$ $=\frac{1}{4N}\sum_{i=1}^{N}\sum_{c_{\pm}=\pm 1}f_{\mathbf{a}}(u_{i,c_{-},c_{+}}(S))$
@@
$=\left(1-\frac{|\mathbf{a}|}{2N}\right)f_{\mathbf{a}}(S)$
$=\lambda_{\mathbf{a}}f_{\mathbf{a}}(S),$ (84)
@@
$\overline{\langle Z_{\mathbf{a}}(t)\rangle}$ $=\sum_{S}p_{t}(S)\bra{\psi_{S}}Z_{\mathbf{a}}\ket{\psi_{S}}=(p_{t}|f_{\mathbf{a}})=(p_{0}|\Gamma^{t}|f_{\mathbf{a}})=\left(1-\frac{|\mathbf{a}|}{2N}\right)^{t}(p_{0}|f_{\mathbf{a}}),$ (85)
@@ D.3 Two particles: relative coordinate eigenmodes
Here we provide more details on eigenmodes of the two-particle problem that depend on subset $S=\{\mathbf{z},\mathbf{z}^{\prime}\}$ only through the relative coordinate $\mathbf{r}=\mathbf{z}\oplus\mathbf{z}^{\prime}$. Let us consider how $\mathbf{r}$ might change under the action of our local automaton gate set $\mathcal{G}=\{u_{iab}\}$. There are two cases to consider:
@@
The rules above describe a random walk on $\mathbb{Z}_{2}^{N}$ based on “activated” bit flips, where $r_{i}$ can flip only if one of its neighbors $r_{i\pm 1}$ is 1. Clearly $\mathbf{r}=\bm{0}$ is a steady state, but it does not describe a valid element of $\Sigma_{2}$ (it gives $\mathbf{z}=\mathbf{z}^{\prime}$). One can check that the restriction to $\mathbb{Z}_{2}^{N}\setminus\{\bm{0}\}$ is irreducible, aperiodic (since lazy moves are allowed) and reversible, thus it has the uniform distribution as unique steady state.
@@
The linear-in-$N$ relaxation time corresponds to the Hamiltonian $H$ being gapped. The Hamiltonian $H$ clearly has two ground states |$\ket{0}^{\otimes N}$ (corresponding to the trivial steady state $\mathbf{r}=\bm{0}$) and |$\ket{+}^{\otimes N}$ (corresponding to the uniform distribution); both states are annihilated by each individual term in $H$, so $H$ is frustration-free. One can use criteria for the existence of gaps in frustration-free, translationally invariant Hamiltonians gosset_local_2016 to establish the presence of a gap as $N\to\infty$ from numerical evaluation of the gap in finite-sized systems. Such an approach reveals that $H$ is indeed gapped as $N\to\infty$ mittal_private_2024 . Note that while the criterion from Ref. gosset_local_2016 is formulated for two-body Hamiltonians, it can be shown to apply for the three-body Hamiltonian $H$ in Eq. (87) because all next-nearest-neighbor terms commute mittal_private_2024 . This establishes the linear-in-$N$ relaxation time for the relative-coordinate Markov chain $\Gamma|_{R}$, consistent with numerical observations in Fig. 6.
@@ Appendix E Properties of the local automaton gate set
In this Appendix we give more details on the local automaton gate set $\mathcal{G}$ used in Sec. 4.3. We recall that the gate set is defined as $\mathcal{G}=\{u_{iab}=\textsf{C}_{i-1,a}\textsf{C}_{i+1,b}{X}_{i}\}$, with $i\in[N]$, $a,b\in\{0,1\}$, and periodic boundary conditions (i.e., $i\pm 1$ are taken modulo $N$). $\textsf{C}_{j,a}$ denotes a control that is activated if bit $j$ is in state $a$. Each time step consists of drawing a gate $u_{iab}\in\mathcal{G}$ uniformly at random and applying it to the system. This defines a random walk on the permutation group $\mathcal{S}_{D}$ with transition matrix
@@
(the dependence on $m$ is left implicit).
@@
It is well known shende_synthesis_2003 ; aaronson_classification_2015 that the Toffoli ($\textsf{CC}X$) and NOT ($X$) gates are “almost” universal for reversible classical computation, in the sense that they generate the subgroup of $\mathcal{S}_{D}$ made of even permutations (permutations that can be decomposed into an even number of transpositions). Our gate set includes Toffoli gates between three consecutive qubits, with the controls on the sides and the target in the middle ($u_{i11}$). Further, it can generate NOT gates ($X_{i}=u_{i00}u_{i01}u_{i10}u_{i11}$) and CNOT gates between neighboring qubits, with either choice of control and target:
@@
Finally, by composing local SWAPs we can make arbitrary permutations of the qubits, and thus generate the whole (non-geometrically-local) Toffoli + NOT gate set. Therefore $\mathcal{G}$ generates all even permutations. Now, given two sets $S,S^{\prime}\in\Sigma_{m}$, we can find a permutation $\sigma\in\mathcal{S}_{D}$ that maps one to the other: $S^{\prime}=\sigma(S)$. If $\sigma$ is even, we are done. If $\sigma$ is odd, then let $i,j\in[D]$ be two basis states that are not in subset $S$ (we have assumed $|S|=m\leq D-2$ so that two such states exist), and let $\tau_{ij}\in\mathcal{S}_{D}$ be the transposition $i\leftrightarrow j$. Then $u\circ\tau_{ij}$ is an even permutation that maps $S$ to $S^{\prime}$. Since any two sets can be connected by an even permutation and any even permutation can be decomposed into a product of $u\in\mathcal{G}$, the Markov chain can connect any two subsets $S,S^{\prime}\in\Sigma_{m}$ and is thus irreducible. ∎
@@
Reversibility, irreducibility and aperiodicity imply that the Markov chain has the uniform distribution $\pi_{m}$ as its unique steady state levin_markov_2017 . This proves that our 3-qubit automaton gate set equilibrates to the uniform distribution on all subset states with $|S|=m\leq D/4$.
It is easy to see that three-qubit gates are necessary to obtain the above result. Indeed two-qubit automaton gates (generated by NOT and CNOT) are affine transformations of the computational basis, in the sense that for any such gate $u$ we have $u\ket{\mathbf{z}}=\ket{A_{u}\mathbf{z}\oplus\mathbf{b}_{u}}$, with $A_{u}$ a binary matrix and $\mathbf{b}_{u}$ a bitstring. Any transformations that are not affine cannot be generated by two-qubit automaton gates. As it turns out, the gate set comprising NOT and CNOT gates induces an irreducible Markov chain only up to $m=3$; an example of two subsets with $m=4$ and $N=3$ that cannot be connected by NOTs and CNOTs only is given by
@@
(each row is a bitstring $\mathbf{z}$ in the set). It is easy to see that the sum (modulo 2) of bitstrings in each set, $\mathbf{z}_{\rm tot}$, transforms under an affine gate $u$ according to $\mathbf{z}_{\rm tot}\mapsto A_{u}\mathbf{z}_{\rm tot}$ (the additive term $\mathbf{b}_{u}$ appears 4 times and thus cancels). But $S$ has $\mathbf{z}_{\rm tot}=(0,0,0)$ while $S^{\prime}$ has $\mathbf{z}_{\rm tot}=(1,1,1)$, so any sequence of $u$’s mapping $S$ to $S^{\prime}$ must be such that $\prod_{u}A_{u}$ annihilates $(1,1,1)$. This is incompatible with a reversible computation (e.g. it would imply a collision between strings $000$ and $111$).
Another way of arriving at a similar conclusion is to note that pseudoentanglement requires magic grewal_stabilizer_2023 ; grewal_pseudoentanglement_2024 , and two-qubit automaton gate sets happen to also be Clifford. Thus if we start from a stabilizer subset state (e.g. |$\ket{0}^{\otimes N-k}\otimes\ket{+}^{\otimes k}$) we cannot produce a pseudoentangled ensemble. However, for $m\leq 3$, it is instead possible to show that the gate set comprising NOT and CNOT yields irreducible Markov chains. This is perhaps related to the fact that Clifford circuits form 3-designs on qubits, so matching the 3 lowest moments of the Haar measure does not require magic.